OpenStudy (anonymous):

Can one person explain to me on how to find the answers? Not just give them, please. Thanks! They are all t/f sin(-136°) = -sin44° tan(-212°) = tan32° sec(-45°) = sec45° cot(-315°) = -cot45° Thanks so much, y'all!

OpenStudy (anonymous):

Hints : \[\sin(-x) = -sinx\] \[\cos(-x) = cosx\] \[\tan(-x) = -tanx\]

OpenStudy (anonymous):

So it's basically just rules I have to memorize? Because my lesson was like subtract from 180, take the square root & multiple by 43. Not really, but it was some long, ridiculous process to find the answer. & It was crazy hard.

OpenStudy (anonymous):

\[Sec(-x)=secx\] \[\cot(-x)=-cotx\]

OpenStudy (anonymous):

sin(-136°) = - sin136 = -sin (180 - 44) = -sin44

OpenStudy (anonymous):

So..1st one is True...Do..Apply Same For All others Keep in mind abt the quadrant in which they Lie

OpenStudy (anonymous):

So I do have to do that thing they taught me? Do I always just subtract from 180? Because there were times they wanted me to take 180 from the angle. Does that make sense?

OpenStudy (anonymous):

& wouldn't that make tan(-212°) = tan32° false since it isn't -tan32°?

hartnn (hartnn):

that is correct, its false. for such problems, u need to know your trigonometric ratios, and its periodicity after how many degrees, value sin repeats itself? and for cos? tan ?

OpenStudy (anonymous):

What do you mean repeat itself? Like start the circle over again?

hartnn (hartnn):

yeah, like sin 90 = 1 and sin (90+360) is also 1 cos 90 = 0 , cos (90+360) is also 0 so yes, drawing a circle would help

OpenStudy (anonymous):

tan(-212°) = - tan(212) = -tan (180 +32) = -tan32

OpenStudy (anonymous):

I have all the values of quadrantal angles if that's what your saying. Like which ones equal 1, 0, -1, and are undefined.

OpenStudy (anonymous):

And when determining which quadrant they are in, I take the number given to me, right? So like -80 would be in the fourth quadrant.

hartnn (hartnn):

yes, -80 is in 4th quad. and in 4th quad, which ratios are negative ?

OpenStudy (anonymous):

sin, tan, csc, cot. (cos & sec are positive)

hartnn (hartnn):

thats correct. u should know that tan values repeat itself after every 180 degrees, and for sin and cos its 360 degrees, then you can make formulas on your own!

OpenStudy (anonymous):

I thought they all repeated every 360 degrees..?

hartnn (hartnn):

sure! tan repeats after 180 means it will repeat after 360 also , right ? and whatever are the properties of sin , cos , tan., it will be same for cosec, sec, cot respectively, so lets talk about sin,cos and tan only

OpenStudy (anonymous):

Okay, I'll try and follow. but I am a tad confused.. Haha (& thanks for helping!)

hartnn (hartnn):

do u want to develop the formula by yourself ? like since tan repeats after 180 , tan (180+x) = tan x tan (x-180) = tan x tan (180-x) = - tan x [because tan -x= -tan x]

OpenStudy (anonymous):

Hm. I would, but I don't know what to do.. & for some reason, those formulas don't make a whole lot of sense to me. I understand how when x & 180 are swapped that it makes your outcome negative. But I don't know how you go to the formula..